Status Report

ISS On-Orbit Status  8 Feb 2003

By SpaceRef Editor
February 8, 2003
Filed under , ,
ISS On-Orbit Status  8 Feb 2003
iss

“We cannot stress enough how blessed and honored we feel to be counted as members of the NASA family. We proudly support the noble goals and objectives of NASA and we will continue to support NASA in its finest and its darkest hours.  It is our deepest hope that you also will continue to share in our belief and support of NASA’s dreams.”   (Public statement [excerpt] by the husbands, wives, and children of the NASA Astronaut Corps, 2/7/03.)

All ISS systems continue to function nominally, except as noted previously or below.  Day 77 in space for the Expedition 6 crew, after a very tough week for all, but especially so for the Columbia crew’s families and also for Ken, Don and Nikolai in their high outpost.

The station crew had their regular off-duty Saturday, except for weekend routine tasks, necessary repair work and daily maintenance.

After wake-up at the regular 1:00am EST, morning inspection, morning hygiene and breakfast, the ISS residents completed the weekly 3-hr. house cleaning.  As part of today’s housekeeping, Bowersox performed head cleaning on VTR-2 (video tape recorder #2), which is getting close to the regular 250-hr. servicing.  [The “uborka stantsii” focuses on removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of surfaces with disinfectants and cleaning of fan screens to avoid temperature rises.]

FE-1 Nikolai Budarin completed the periodic inspection of the BRPK air/condensate separator of the SRVK water processing system.   [The BRPK separator contains porous cermet hydrophilic (“water attracting”) tubes through which the gas-liquid mixture from the heat exchanger moves.  They separate the air from the condensate, but when the separator exceeds its service life or malfunctions, incomplete separation of the atmospheric condensate occurs, and the water then collects under the “sheet” of porous fluoroplastic.  This is the main focus of the regular inspection.]

Budarin, supported by FE-2/SO Don Pettit, then removed and replaced a pipe conduit at the BPK condensate micropump (MR2-2G) of the BRPK separator unit.  [The pipe conduit connects the BPK MR2-2G to a solenoid valve on the BPK air line.  Objective of the 2-hr. installation of the new conduit, which is fitted with a check valve and control valves, is to protect the micropump from moisture ingestion and decrease the flow of supply air, thereby mitigating the adverse effect that the operation of the BPK unit has had on the life span of the BRPK air/condensate separators.]

On MCC-M Go-ahead, Budarin also terminated the regeneration cycle on adsorbent bed #1 of the BMP harmful impurities filter unit, returning it to Purify mode.  Later, after #1 was determined to work properly, he switched channel #2 to the 24-hr bake-out process for regeneration.

At 4:00am and 5:50am, FE-1 Budarin conducted two downlink sessions of the video records of the TEKh-20 Plasma Crystal experiment (PK-3) conducted on 1/22.  [After playback of the videocassette, Nikolai left all Russian Telescience components connected for the next downlinks, currently scheduled for 2/12 (next Wednesday).]

Nikolai also performed the task-list item of reviewing his upcoming next session with the MBI-8 biomedical “Profilaktika” (preventive health maintenance) fitness session, scheduled for 2/10-12.  [During these three days, Budarin will use the VELO ergometer on Day 1, the Load Trainer on Day 2 and the TVIS treadmill on Day 3. The test is similar to the MedOps MO-3 test, except that it uses the TEEM-100M gas analyzer during exercise operations, blood lactate determination and subjective load assessment.]

Don Pettit again collected the two regular daily carbon dioxide (CO2) readings in the SM and Lab with the U.S. CDMK (CO2 monitoring kit).

The daily routine inspection/maintenance of the SOSh life support system, incl. the ASU toilet system, was completed by CDR Ken Bowersox, while Pettit prepared the IMS inventory system for database auto import/export. [After database update files are uplinked, the crew needs to auto import them into the system].

Budarin conducted his daily 24-hr. checkup and watering of the Russian BIO-5 Rasteniya-2/Lada-2 (“Plants-2”) plant growth experiment.

For “Diatomeya” world ocean observations, on the Russian task list, Budarin today was to observe and image seawater bloom features (TsKO) and cloud structures in the waters of the Atlantic, Pacific, and Indian Oceans.  [Bioproductive areas were identified in the Atlantic west of South Africa, in the Benguela Current and Upwelling and along the south boundary of the Gulf Stream, in the Pacific in the frontal portion of the Peruvian Current.]
 
Today’s targets for the Russian “Uragan” (hurricane) earth observation experiment, also a task list item using the Nikon D1 digital still camera (f80-400 zoom lens), were the Himalayas (convergence shooting, i.e., a series of consecutive frames while keeping the target at the center of the camera’s field of view), the city of Burgas and the Panama Canal.

At 7:05am EST, attitude control was handed over to the Russian MCS (motion control system) thrusters of the Service Module for the dynamic test of the US-21 matching unit installed in the Progress 10P yesterday.  [The test firings of Progress thruster manifolds 1 and 2 were to be completed at 10:48am, and control authority handed back to U.S. CMGs (control moment thrusters) at 11:00am to resume momentum management in LVLH (local vertical/local horizontal).  Progress thruster yaw and pitch control was incorporated into the SM’s propulsion control algorithm.]

PFCs (private family conferences) were scheduled for Ken Bowersox and Don Pettit today and are timelined for all crewmembers tomorrow.

Yesterday’s Velocicalc airflow measurements determined low air flow from the Node aft port and Lab aft port fans.  [To support troubleshooting, the crew was asked tocheckthe position of the two Node air damper valves.  Also, the ground will deactivate and reactivate each suspect fan to determine the run and startupcharacteristics. As a comparison, the starboard aft fan will be likewisebe deactivated and activated.Further troubleshooting will follow next week but as of now there is no other action required on the crew’s part.]

MSG (microgravity science glovebox) continues inoperative.  Troubleshooting (t/s) efforts are underway at POC (Payload Operations Center) by ESA.  [In order to minimize the risk of additional problems associated with closing a circuit breaker into a possible shorted circuit, the t/s will wait until ESA and MSFC have completed a detailed audit of the as-built MSG configuration, analyzed the failure data, and re-reviewed ground test data.  Initial t/s procedures are expected to be available in the next few days, and some time has been scheduled on the crew’s timeline early next week to allow direct discussion with the MSG ground team.  A detailed information package on MSG activities has also been uplinked.]

Science Update (Expedition Six — 10th):

The crew was thanked by the Lead Increment Scientist for attempting the MSG repair and collecting data for the PuFF, EVARM, and Foot investigations this week.

Extra-Vehicular Activity Radiation Monitors (EVARM):   Downlinked data will provide excellent insight into the interior of ISS.

GASMAP/Pulmonary Function in Flight (PuFF):  The ground had Ku-band reception for about two-thirds of each session this week, and the real-time downlink confirmed what was already known: “the crew knows how to breathe”.  PuFF support teams at both UCSD (University of California, San Diego) and JSC watch every breath the crew takes, along with every calibration, whenever they have Ku coverage.  “There are always cheers when the cal numbers are close and the MEFV blows are vertical.”
 
Renal (Kidney) Stone Experiment:  Completed for Increment 6.  Bowersox and Pettit continue to take pills.

Human Research Facility/Workstation (HRF WS):   Continuing.

Foot/Ground Reaction Forces During Space Flight (FOOT):  Good run by Bowersox this week.

Pore Formation and Mobility Investigation (PFMI):  Complete for Inc. 5.  On hold pending MSG PDC (power distribution controller) and ESEM3 (exchangeable standard electronic module 3) replacement. Will remain in MSG until further notice.

Space Acceleration Measurement System (SAMS):  SAMS continues to collect acceleration data for vibratory characterization of the microgravity environment. Both SAMS and MAMS captured Progress 10P docking earlier this week.  Both instruments will measure the accelerations during the reboost activities planned for next week.

Microgravity Acceleration Measurement System (MAMS):  MAMS continues measurement of microgravity environment in the quasi-steady regime for general characterization. 

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):   Temperatures are nominal. PCG-STES is currently in heating mode.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):   On hold until the MSG (microgravity science glovebox) is operational.

Materials ISS Experiment (MISSE):  In progress.  Deployed outside. Nominal and collecting data.  Was photographed by the crew during the 1/15 EVA.

Zeolite Crystal Growth (ZCG):  ZCG has finished science operations for Inc 6.

EarthKAM (EK):  All activities complete for Increment 6.

Crew Earth Observation (CEO):  The quality and composition of CEO downlink imagery continues to improve.  One of the crew’s dramatic views of bushfires in southeastern Australia was selected for Earth Observatory publication this week.  The rate of review of ISS/CEO imagery will be slower now due to staff reassignment.
Today’s targets for the CEO (crew earth observations) program were Beijing, China (with fair weather in winter sun, looking just left of track for views of the Chinese capital city near the base of the Luliang Range), Tianjin, China (looking for nadir views of this major port city for northern China, near the coast about 60 miles southeast of Beijing), Industrialized Southeastern Africa (a high-pressure area bearing aerosols from the industrialized interior of South Africa is slowly moving off the SE coast. On this pass the crew was to try oblique views to detect the extent of the aerosol plume as it exits the coast over the darker waters, and to include some coastline in their views for reference points), Eastern Mediterranean Dust (satellite imagery continues to show dust blowing out to sea from the northern coast of Egypt in response to a winter storm over western Turkey. As ISS tracked northeastward over the Sinai Peninsula, crew was to take images left of track towards the Mediterranean in oblique views), Eastern Mediterranean Smog (a stable high-pressure area has settled over the central Mediterranean. As ISS approached the coast of Greece, the crew was to shoot obliquely to the left of track for aerosol plumes moving down the Adriatic Sea and off the coast of southern Italy), Western Mediterranean Smog (on this pass the crew was to look for indications of aerosols exit the coast of northeastern Spain and southern France, especially from the Rhone River valley. As ISS crossed the E coast of Spain, they were to look left of track), Tuamotu-Austral Islands (crew was to use the long lenses of the digital camera to obtain detailed nadir views of the small reefs and atolls of this archipelago), and Tuamotu Archipelago (ISS pass was across a broad and dense segment of this large island chain.  Long lens digital images at nadir will be used by international researchers to map and inventory the reefs).
CEO images can be viewed at the website
http://eol.jsc.nasa.gov

ISS Orbit  (as of this morning, 7:06am EST [= epoch]):

  • Mean altitude — 386.1 km
  • Apogee — 390.0 km
  • Perigee — 382.3 km
  • Period — 92.28 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0005674
  • Orbits per 24-hr. day — 15.60
  • Mean altitude loss in last 24 hours — 175 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 24092
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see
  •  http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.